# Staying on the Manifold: Geometry-Aware Noise Injection

### v1

@@ Abstract
It has been shown that perturbing the input during training implicitly regularises the gradient of the learnt function, leading to smoother models and enhancing generalisation. However, previous research mostly considered the addition of ambient noise in the input space, without considering the underlying structure of the data. In this work, we propose several methods of adding geometry-aware input noise that accounts for the lower dimensional manifold the input space inhabits. We start by projecting ambient Gaussian noise onto the tangent space of the manifold. In a second step, the noise sample is mapped on the manifold via the associated geodesic curve. We also consider Brownian motion noise, which moves in random steps along the manifold. We show that geometry-aware noise leads to improved generalization and robustness to hyperparameter selection on highly curved manifolds, while performing at least as well as training without noise on simpler manifolds. Our proposed framework extends to learned data manifolds.
@@ 1 Introduction
One of the most intuitive and practical methods to improve the generalisation properties of a learnable model is to consider data augmentation techniques [1]. During training, new data samples are created from given ones, sharing the same features and labels. This approach has been extensively used with image data, for example through adjusting the illumination, changing the orientation or cropping.
@@
Classic machine learning research has already established the influence of input noise on generalisation performance [2, 3]. One widely studied technique is adding Gaussian noise to the inputs, which leads to a smoothness penalty on the learnt function [4, 5], however, these works do not take into account the structure of the input data. A fundamental observation in modern machine learning is the manifold hypothesis: it states that high-dimensional data tends to concentrate around a lower-dimensional manifold in the ambient space [7, 6]. In the context of noise-based learning, this has the implication that, with high probability, Gaussian noise will be almost perpendicular to the manifold [8]. Hence, adding Gaussian noise to the data leads to unlikely or non-informative augmented data samples.
@@
Figure 1: Noise injection is a data augmentation technique that can improve generalisation. For a data point ( ) lying on a lower-dimensional manifold, sampling noise in the ambient space ( ) almost surely deviates from the input manifold whereas a sample from a geometry-aware noise process ( ) stays on the manifold and respects the data geometry. Illustrated of the biconcave disc that resembles a red blood cell.
@@
Additionally, many real-world problems require learning functions on a known manifold rather than the unconstrained Euclidean space. Weather and climate observations naturally live on the surface of the sphere, which approximates the shape of the Earth. In cell biology we might consider red blood cells, which can be approximated by a biconcave disc [9]. Or in brain imaging, quantities like cortical thickness and grey matter intensity are measured on the cortical surface [10]: although the cortex can be mapped onto the sphere, it is actually highly wrinkly. In such settings, applying perturbations or learning representations that ignore the intrinsic manifold structure can lead to deceptive results as Euclidean distances in the embedding space fail to capture the true distances between points: two points which might be close with respect to the Euclidean metric can be far apart when travelling along the manifold surface. This highlight the necessity of geometry-aware methods that respect the manifold structure when perturbing data as an augmentation technique.
@@
In this paper, we propose geometry-aware noise injection strategies as a data augmentation technique and show their benefits compared to ambient space noise injection. We consider three such strategies – tangent space noise, geodesic noise and Brownian motion noise – and demonstrate their effect on manifolds embedded in $\mathbb{R}^{3}$, namely the Swiss roll and families of spheroids and tori.
Our contributions include:
  • definitions, derivations, and implementations of geometry-aware input noise for various parametrised manifolds and their deformations,
@@
  • empirical demonstration that geometry-aware noise improves generalisation and robustness over manifold-agnostic noise.
@@ 2 Preliminaries
We consider a dataset of $N$ points $\{{\boldsymbol{x}}_{n},y_{n}\}_{n=1}^{N}$, where the inputs ${\boldsymbol{x}}_{n}\in\mathcal{X}=\mathbb{R}^{D}$ are assumed to lie on an embedded $d$-dimensional manifold $\mathcal{M}$ with $d<D$, and the outputs $y_{n}\in\mathcal{Y}$ may be either continuous or discrete. Our goal is to learn a function $f_{{\boldsymbol{\theta}}}:\mathcal{X}\rightarrow\mathcal{Y}$, typically parametrised by a deep neural network with parameters ${\boldsymbol{\theta}}\in\mathbb{R}^{K}$. The model is trained by minimizing the empirical loss
$\mathcal{L}({\boldsymbol{x}},{\boldsymbol{\theta}})=\sum_{n=1}^{N}\ell(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}),y_{n}),$
@@ 2.1 Gaussian Input Noise
As mentioned, several previous works consider Gaussian input noise [4, 11, 12, 2]. In this section, we summarise the previous analysis and show that adding Gaussian noise to the input during training is equivalent in expectation to Tikhonov regularisation [13].
Consider an input data point ${\boldsymbol{x}}_{n}\in\mathcal{X},$ which we perturb with noise following a normal distribution ${\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{D})$ for $\sigma>0$. Then the second-order Taylor expansion of the loss function $\mathcal{L}\left({\boldsymbol{x}}\right)$ is given by:
@@
We take the expectation of the Gaussian noise distribution and get
@@
where $\Delta$ is the Laplace operator (trace of the Hessian). Using the chain rule, this term expands to:
$\Delta_{x}\mathcal{L}\left({\boldsymbol{x}}\right)$ $=\sum_{n=1}^{N}\lVert\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\rVert^{2}$ (3)
$+\frac{1}{2}\sum_{n=1}^{N}\left(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})-y_{n}\right)\Delta_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}).$
when choosing $\ell$ to be the MSE. When the function interpolates the training data points, that is, $f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\approx y_{n}$, the second summand in Equation 3 vanishes. Thus, after plugging this back into Equation 2, we see that adding input noise is equivalent (in expectation) to optimising a regularised loss on the form $\mathcal{L}\left({\boldsymbol{x}}\right)+R\left({\boldsymbol{x}},{\boldsymbol{\theta}}\right),$ with $R$ being the Tikohonov regulariser
$R\left({\boldsymbol{x}},{\boldsymbol{\theta}}\right)=\frac{\sigma^{2}}{2}\sum_{n=1}^{N}\lVert\nabla_{x}f_{\theta}({\boldsymbol{x}}_{n})\rVert^{2}.$ (4)
@@ 2.2 Riemannian Geometry
In this section, we give a brief introduction to the tools we use from Riemannian geometry [14]. The reader who is already familiar with Riemannian geometry may skip this section.
@@
A manifold $\mathcal{M}$ is a Hausdorff space such that for every $\boldsymbol{x}$$\in\mathcal{M}$ there exists a homeomorphism $X:U\rightarrow V$ from a neighbourhood $U\ni{\boldsymbol{x}}$ to an open set $V\subseteq\mathbb{R}^{d}$. Further, we require these charts to be compatible on the intersection of their domains, i.e.
@@
In $\mathbb{R}^{3},$ the tangent plane of a manifold is easy to picture: each point of the surface is approximated with a plane in which the tangent vectors live. In higher dimensions, we say that the tangent space $T_{{\boldsymbol{x}}}\mathcal{M}$ of $\mathcal{M}$ at a point $\boldsymbol{x}$} consists of the velocities of all curves on $\mathcal{M}$ passing through $\boldsymbol{x}$}, that is, if $\gamma$ is a smooth curve on $\mathcal{M}$ parametrised by time $t$ with
$\gamma(0)={\boldsymbol{x}},$
then
$\boldsymbol{v}=\dot{\gamma}(0)\in T_{{\boldsymbol{x}}}\mathcal{M}.$
Assume we have a smooth parametrisation $X:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}.$ Then the Jacobian
$\mathbf{J}_{X}=\left[\frac{\partial X}{\partial u_{1}},\dots,\frac{\partial X}{\partial u_{d}}\right]$
is a function from $\mathbb{R}^{d}$ to $\mathbb{R}^{D\times d}$ and the tangent space at each point is spanned by the columns of $\mathbf{J}_{X}.$ At every point $\boldsymbol{x}$$\in\mathcal{M}$, any vector $\boldsymbol{v}\in\mathbb{R}^{D}$ can be orthogonally decomposed into a tangential and a normal component as
$\boldsymbol{v}=\boldsymbol{v}_{\top}+\boldsymbol{v}_{\perp}.$
In Figure 2, we can see a manifold (the sphere) embedded in $\mathbb{R}^{3},$ and the tangent space at a point.
@@
A Riemannian manifold $\left(\mathcal{M},g\right)$ is a smooth manifold equipped with a Riemannian metric. A metric $g$ of $\mathcal{M}$ equips each point $\boldsymbol{x}$$\in\mathcal{M}$ with an inner product $g_{{\boldsymbol{x}}}$ on $T_{{\boldsymbol{x}}}\mathcal{M}.$ This continuous tensor field allows us to measure distances and angles, and define geodesics along the manifold.
Assume we have a smooth parametrisation $X:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}.$ Then the matrix valued function
$\mathbf{J}_{X}^{\top}\cdot\mathbf{J}_{X}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2\times 2}$
induces a metric. For $X(\boldsymbol{u})={\boldsymbol{x}}\in\mathcal{M}$ and $\boldsymbol{v},\boldsymbol{w}\in T_{{\boldsymbol{x}}}\mathcal{M},$ let $\tilde{\boldsymbol{v}},\tilde{\boldsymbol{w}}\in T_{\boldsymbol{u}}\mathbb{R}^{d}$ be such that $\mathbf{J}_{X}\tilde{\boldsymbol{v}}=\boldsymbol{v}$ and $\mathbf{J}_{X}\tilde{\boldsymbol{w}}=\boldsymbol{w}$. Then
$g_{{\boldsymbol{x}}}(\boldsymbol{v},\boldsymbol{w})=\boldsymbol{v}^{\top}\mathbf{J}_{X}^{\top}\mathbf{J}_{X}\boldsymbol{w}$
is the induced metric. By abuse of notation, we will often write $g$ to denote the matrix $\mathbf{J}_{X}^{\top}\mathbf{J}_{X}$.
@@
A geodesic is locally the shortest path on a manifold. We can rewrite a curve
$\gamma:I\subseteq\mathbb{R}\rightarrow\mathcal{M}$
on $\mathcal{M}$ as $\gamma(t)=X\circ\alpha(t),$ where $\alpha:I\rightarrow\mathbb{R}^{d}$ is a curve in the parameter space. Then $\gamma$ is a geodesic if and only if $\alpha$ satisfies the following ordinary differential equation (ODE) for all $k=1,\dots,d$:
@@
$\gamma_{\boldsymbol{e}}(0)={\boldsymbol{x}},\quad\dot{\gamma}_{\boldsymbol{e}}(0)=\boldsymbol{e}.$
@@
Intuitively, one can imagine the exponential map like a function which wraps aluminium foil (the tangent plane) around, say, a bagel (the manifold). Despite the manifold being curved and the tangent space being flat, at any chosen point we can wrap a small part of the tangent plane around a neighbourhood of this point without folding it.
Using geodesics, for each $\boldsymbol{x}$$\in\mathcal{M}$ we can define a map from an open ball $B_{\delta}(0)\subseteq T_{{\boldsymbol{x}}}\mathcal{M}$ to a neighbourhood ${\boldsymbol{x}}\in U\subseteq\mathcal{M}$ on the manifold. We will call this map the exponential map: [^1] Here, $\delta\in\mathbb{R}^{+}$ ensures that the exponential map is a well defined diffeomorphism. Loosely speaking, it is the largest radius we can choose while guaranteeing that the geodesics are well defined and do not overlap.
$\operatorname{Exp}_{{\boldsymbol{x}}}:B_{\delta}(0)\subseteq T_{{\boldsymbol{x}}}\mathcal{M}\rightarrow U\subseteq\mathcal{M},$
@@
In other words, the exponential map maps a vector $\boldsymbol{v}\in T_{{\boldsymbol{x}}}\mathcal{M}$ from the tangent space to the endpoint of a curve on the manifold, $\gamma_{\frac{v}{\lVert\boldsymbol{v}\rVert}}(\lVert\boldsymbol{v}\rVert),$ and the zero vector to $\boldsymbol{x}$}.

### v2

@@ Abstract
It has been shown that perturbing the input during training implicitly regularises the gradient of the learnt function, leading to smoother models and enhancing generalisation. However, previous research mostly considered the addition of ambient noise in the input space, without considering the underlying structure of the data. In this work, we propose several strategies of adding geometry-aware input noise that accounts for the lower dimensional manifold the input space inhabits. We start by projecting ambient Gaussian noise onto the tangent space of the manifold. In a second step, the noise sample is mapped on the manifold via the associated geodesic curve. We also consider Brownian motion noise, which moves in random steps along the manifold. We show that geometry-aware noise leads to improved generalisation and robustness to hyperparameter selection on highly curved manifolds, while performing at least as well as training without noise on simpler manifolds. Our proposed framework extends to data manifolds approximated by generative models and we observe similar trends on the MNIST digits dataset.
@@ 1 Introduction
One of the most intuitive and practical methods to improve the generalisation properties of a learnable model is to consider data augmentation techniques [lecun2002gradient]. During training, new data samples are created from given ones, sharing the same features and labels. This approach has been extensively used with e.g. image data, through adjusting the illumination, changing the orientation or cropping.
@@
Classic machine learning research has already established the influence of input noise on generalisation performance [sietsma1991creating, ferianc2024navigatingnoisestudynoise]. One widely studied technique is adding Gaussian noise to the inputs, which leads to a smoothness penalty on the learnt function [10.1162/neco.1995.7.1.108, an1996effects], however, these works do not take into account the structure of the input data. A fundamental observation in machine learning is the manifold hypothesis: it states that high-dimensional data tends to concentrate around a lower-dimensional manifold in the ambient space [fefferman2016testing, belkin2001laplacian]. In the context of noise-based learning, this has the implication that, with high probability, Gaussian noise will be almost perpendicular to the manifold [fefferman2023fitting]. Hence, Gaussian input noise gives unlikely or non-informative augmented data samples.
@@
Figure 1: Noise injection is a data augmentation technique that can improve generalisation. For a data point ( ) lying on a lower-dimensional manifold, sampling noise in the ambient space ( ) almost surely deviates from the input manifold whereas a sample from a geometry-aware noise process ( ) stays on the manifold and respects the data geometry. Illustration of the biconcave disc that resembles a red blood cell.
@@
Additionally, many real-world problems require learning functions on a known manifold rather than the unconstrained Euclidean space. Weather and climate observations naturally live on the surface of the sphere, which approximates the shape of the Earth. In cell biology we might consider red blood cells, which can be approximated by a biconcave disc [kuchel2021surface]. Or in brain imaging, quantities like cortical thickness and grey matter intensity are measured on the cortical surface [ripart]: although the cortex can be mapped onto the sphere, it is actually highly wrinkly. In such settings, applying perturbations or learning representations that ignore the intrinsic manifold structure can lead to deceptive results as Euclidean distances in the embedding space fail to capture the true distances between points: two points which might be close with respect to the Euclidean metric can be far apart when travelling along the manifold surface. This highlights the necessity of geometry-aware strategies that respect the manifold structure when perturbing data as an augmentation technique.
@@
In this paper, we propose geometry-aware noise injection strategies as a data augmentation technique and show their benefits compared to ambient space noise injection. We consider three such strategies and demonstrate their effect on manifolds embedded in $\mathbb{R}^{3}$, namely the Swiss roll and families of spheroids and tori. We additionally apply our strategies in the setting of a learnt data manifold, specifically the MNIST digits dataset. Our contributions include:
  1. 1.
    defining geometry-aware input noise for various parameterised, deformed and learned manifolds,
  2. 2.
@@
  3. 3.
    empirical demonstration that geometry-aware noise can improve generalisation and robustness over manifold-agnostic noise.
@@ 2 Preliminaries
We consider a dataset of $N$ points $\{{\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n}\}_{n=1}^{N}$, where the inputs ${\boldsymbol{x}}_{n}\in\mathcal{X}\subseteq\mathbb{R}^{D}$ are assumed to lie on an embedded $d$-dimensional manifold $\mathcal{M}$ with $d<D$, and the outputs ${\boldsymbol{y}}_{n}\in\mathcal{Y}$ may be either continuous or discrete. Our goal is to learn a function $f_{{\boldsymbol{\theta}}}:\mathcal{X}\rightarrow\mathcal{Y}$, typically parameterised by a deep neural network with parameters ${\boldsymbol{\theta}}\in\mathbb{R}^{K}$. The model is trained by minimizing the empirical loss
$\mathcal{L}({\boldsymbol{x}},{\boldsymbol{\theta}})=\sum_{n=1}^{N}\ell(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}),{\boldsymbol{y}}_{n}),$ (1)
@@ 2.1 Gaussian Input Noise
Several previous works consider Gaussian input noise [10.1162/neco.1995.7.1.108, matsuoka1992noise, rifai2011adding, sietsma1991creating]. In this section, we summarise the previous analysis and show that adding Gaussian noise to the input during training is equivalent in expectation to Tikhonov regularisation [tikhonov1977solutions].
Consider an input data point ${\boldsymbol{x}}_{n}\in\mathcal{X},$ which we perturb with noise following a Gaussian distribution ${\boldsymbol{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}\mathbb{I}_{D})$ for $\sigma>0$. Then the second-order Taylor expansion of the loss function $\mathcal{L}\left({\boldsymbol{x}}\right)$ is:
@@
Taking the expectation of the Gaussian noise yields
@@
where $\Delta_{x}$ is the Laplace operator (trace of the Hessian) with respect to $x$. When choosing $\ell$ to be the MSE, and using the chain rule, this expands to:
$\Delta_{x}\mathcal{L}\left({\boldsymbol{x}}\right)$ $=\frac{1}{N}\cdot\sum_{n=1}^{N}\lVert\nabla_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\rVert^{2}$ (4)
$+\frac{1}{2N}\sum_{n=1}^{N}\left(f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})-{\boldsymbol{y}}_{n}\right)\Delta_{x}f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n}).$
When the function interpolates the training data points, that is, $f_{{\boldsymbol{\theta}}}({\boldsymbol{x}}_{n})\approx{\boldsymbol{y}}_{n}$, the second summand in Equation 4 vanishes [^1] We assume that $\Delta_{x}f_{\theta}({\boldsymbol{x}}_{n})$ is bounded for all ${\boldsymbol{x}}_{n}$ and $\theta,$ as $\mathcal{C}^{2}$-smoothness is satisfied globally for several activation functions, e.g. Softplus and Tanh. Since the set of training points is finite, we conclude that $\Delta_{x}f_{\theta}({\boldsymbol{x}}_{n})$ is bounded. For ReLU architectures, which are not $\mathcal{C}^{2}$-smooth, the set of cusps has measure zero, and $\nabla_{x}^{2}f_{\theta}({\boldsymbol{x}})$ vanishes almost everywhere.. Thus, after plugging this back into Equation 3, we see that adding input noise is equivalent (in expectation) to optimising a regularised loss on the form $\mathcal{L}\left({\boldsymbol{x}}\right)+R\left({\boldsymbol{x}},{\boldsymbol{\theta}}\right),$ with $R$ being the Tikhonov regulariser
$R\left({\boldsymbol{x}},{\boldsymbol{\theta}}\right)=\frac{\sigma^{2}}{2N}\sum_{n=1}^{N}\lVert\nabla_{x}f_{\theta}({\boldsymbol{x}}_{n})\rVert^{2}.$ (5)
@@ 2.2 Riemannian Geometry
A manifold $\mathcal{M}$ is a Hausdorff space such that for every $\boldsymbol{x}$$\in\mathcal{M}$ there exists a homeomorphism $X:U\rightarrow V$ from a neighbourhood $U\ni{\boldsymbol{x}}$ to an open set $V\subseteq\mathbb{R}^{d}$. We require these charts to be compatible on the intersection of their domains, i.e.
@@
In $\mathbb{R}^{3},$ the tangent plane of a manifold is easy to picture: each point of the surface is approximated with a plane in which the tangent vectors live. In higher dimensions, we say that the tangent space $T_{{\boldsymbol{x}}}\mathcal{M}$ of $\mathcal{M}$ at a point $\boldsymbol{x}$} consists of the velocities of all curves on $\mathcal{M}$ passing through $\boldsymbol{x}$}, that is, if $\gamma$ is a smooth curve on $\mathcal{M}$ parameterised by time $t$ with $\gamma(0)={\boldsymbol{x}},$ then $\boldsymbol{v}=\dot{\gamma}(0)\in T_{{\boldsymbol{x}}}\mathcal{M}.$ Assume we have a smooth parameterisation $X:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}.$ Then the Jacobian of the chart,
$\mathbf{J}_{X}=\left[\frac{\partial X}{\partial u_{1}},\dots,\frac{\partial X}{\partial u_{d}}\right]$ (6)
is a function from $\mathbb{R}^{d}$ to $\mathbb{R}^{D\times d}$ and the tangent space at each point is spanned by the columns of $\mathbf{J}_{X}.$ At every point $\boldsymbol{x}$$\in\mathcal{M}$, any vector $\boldsymbol{v}\in\mathbb{R}^{D}$ can be orthogonally decomposed into a tangential and a normal component as $\boldsymbol{v}=\boldsymbol{v}_{\top}+\boldsymbol{v}_{\perp}.$ In Figure 2, we show a manifold (the sphere) embedded in $\mathbb{R}^{3},$ and the tangent space at a point.
@@
A Riemannian manifold $\left(\mathcal{M},g\right)$ is a smooth manifold equipped with a Riemannian metric. A metric $g$ of $\mathcal{M}$ equips each point $\boldsymbol{x}$$\in\mathcal{M}$ with an inner product $g_{{\boldsymbol{x}}}$ on $T_{{\boldsymbol{x}}}\mathcal{M}.$ This tensor field allows us to measure distances and angles on the manifold. Given a smooth parameterisation $X:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D},$ the matrix valued function
$\mathbf{J}_{X}^{\top}\cdot\mathbf{J}_{X}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ (7)
induces a metric. For $X(\boldsymbol{u})={\boldsymbol{x}}\in\mathcal{M}$ and $\boldsymbol{v},\boldsymbol{w}\in T_{{\boldsymbol{x}}}\mathcal{M},$ let $\tilde{\boldsymbol{v}},\tilde{\boldsymbol{w}}\in T_{\boldsymbol{u}}\mathbb{R}^{d}$ be such that $\mathbf{J}_{X}\tilde{\boldsymbol{v}}=\boldsymbol{v}$ and $\mathbf{J}_{X}\tilde{\boldsymbol{w}}=\boldsymbol{w}$. Then the induced metric is
$g_{{\boldsymbol{x}}}(\boldsymbol{v},\boldsymbol{w})=\boldsymbol{v}^{\top}\mathbf{J}_{X}^{\top}\mathbf{J}_{X}\boldsymbol{w}.$ (8)
We will often write $g$ to denote the matrix $\mathbf{J}_{X}^{\top}\mathbf{J}_{X}$.
@@
A geodesic is locally the shortest path on a manifold. We can write a curve $\gamma:I\subseteq\mathbb{R}\rightarrow\mathcal{M}$ on $\mathcal{M}$ as $\gamma(t)=X\circ\alpha(t),$ where $\alpha:I\rightarrow\mathbb{R}^{d}$ is a curve in the parameter space. Then $\gamma$ is a geodesic if and only if $\alpha$ satisfies the following ordinary differential equation (ODE) for all $k=1,\dots,d$:
@@
$\gamma_{\boldsymbol{e}}(0)={\boldsymbol{x}},\quad\dot{\gamma}_{\boldsymbol{e}}(0)=\boldsymbol{e}.$ (10)
@@
One can imagine the exponential map as a function which wraps aluminium foil (the tangent plane) around some object (the manifold). Though the manifold is curved and the tangent space is flat, we can wrap a small part of the tangent plane around a neighbourhood of any point without folding the plane.
Using geodesics, for each $\boldsymbol{x}$$\in\mathcal{M}$ we can define a map from an open ball $B_{\delta}(0)\subseteq T_{{\boldsymbol{x}}}\mathcal{M}$ of radius $\delta$ to a neighbourhood ${\boldsymbol{x}}\in U\subseteq\mathcal{M}$ on the manifold [^2] Here, $\delta\in\mathbb{R}^{+}$ ensures that the exponential map is a well defined diffeomorphism. Loosely speaking, it is the largest radius we can choose while guaranteeing that the geodesics are well defined and do not overlap., i.e. $\operatorname{Exp}_{{\boldsymbol{x}}}:B_{\delta}(0)\subseteq T_{{\boldsymbol{x}}}\mathcal{M}\rightarrow U\subseteq\mathcal{M}$. We will call this map the exponential map and define it as:
@@
Hence, the exponential map maps a tangent space vector $\boldsymbol{v}\in T_{{\boldsymbol{x}}}\mathcal{M}$ to the endpoint of a curve on the manifold, $\gamma_{\frac{v}{\lVert\boldsymbol{v}\rVert}}(\lVert\boldsymbol{v}\rVert),$ and the zero vector to $\boldsymbol{x}$}.